OpenStudy (jazzyluvsu123):

ok now I just need answers cause I gotta get to bed The teacher to Freshman student ratio at the high school is 1 to 27. if there are 810 freshman students, how many teachers are needed? A. 31 B. 27 C. 30 D. 29

OpenStudy (anonymous):

\[\frac{27}{1}=\frac{x}{8100}\iff x=\frac{8100}{27}\]

OpenStudy (jazzyluvsu123):

u mean 810 ? @satellite73

OpenStudy (jazzyluvsu123):

is my answer 30 @iambatman